Two cars A and B take part in a race 700 km long.
Car A travels an average of 35 km per 12 minutes.
Whereas Car B travels at an average of 58.3 km per 20 minutes.
by what distance does the faster car beat the other car in the race?

Answer by mananth(16946)

Two cars A and B take part in a race 700 km long.
disance =700
Car A travels an average of 35 km per 12 minutes.
35*60/12 = 175 km/h
time taken = 700/175 =4 hours
Whereas Car B travels at an average of 58.3 km per 20 minutes.
58.3*60/20 = 174.9 km/h
time taken = 700/174,9
=4.0022 hours
in 0.0022 hours Car A travels 175*0.0022
=0.385 km
beats by 385m